OpenStudy (anonymous):

2. Simplify the given expression to rational exponent form and justify each step by identifying the properties of rational exponents used. All work must be shown. 1/ (3 root x^-6)

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@hartnn

OpenStudy (phi):

is this \[ \frac{1}{\sqrt[3]{x^{-6} }} \]

OpenStudy (anonymous):

yes

OpenStudy (phi):

you need to remember 3 things: 1) when you see "root" think EXPONENT IS A FRACTION example: square root means exponent 1 /2 cube root means exponent is 1 /3 n th root means exponent is 1 / n can you use that idea to rewrite \[ \sqrt[3]{x^{-6}} \] ? put parens around the "stuff" inside the radical sign, and put an exponent 1 /3 to the upper right of the parens

OpenStudy (anonymous):

\[(\sqrt{-18x})^3\]

OpenStudy (phi):

don't change what is inside the radical sign (the \( \sqrt{} \) sign) in other words \(x^{-6}\) does not change into \( -18x\) also, read carefully: 1) REMOVE the radical sign. 2) put parens around what's left 3) make the exponent 1 / 3 (notice 3rd root becomes exponent 1 /3

OpenStudy (anonymous):

\[(x^{-6})^3\]

OpenStudy (phi):

almost. but the exponent should be 1 /3 (that means the 3rd root)

OpenStudy (anonymous):

so (x^-6)^1/3

OpenStudy (phi):

yes. The important "rule" is change the radical to a FRACTIONAL exponent. next we use rule #2 \[ \left(a^b\right)^c = a^{bc} \] we MULTIPLY the exponents to get rid of the parens can you do that ?

OpenStudy (anonymous):

(x^-6)^1/3 x^-2 since -6 times 1/3 is 2

OpenStudy (phi):

yes, but I would say since -6 times 1/3 is -2

OpenStudy (phi):

so far we have \[ \frac{1}{\sqrt[3]{x^{-6} }} = \frac{1}{x^{-2} }\] rule #3: if we "flip" a fraction, we change the sign of the exponent: \[ \frac{1}{a^{-b}} = a^b \] or \[ \frac{1}{a^{b}} = a^{-b} \] can you use that rule on \[ \frac{1}{x^{-2} }\]?

OpenStudy (anonymous):

\[\frac{ x^2 }{ 1 }\]

OpenStudy (phi):

yes, and anything divided by 1 is itself in other words x^2 / 1 = x^2

OpenStudy (anonymous):

so the answer is x^2

OpenStudy (phi):

yes. Hopefully if you learn the 3 rules, you can do any of these problems.

OpenStudy (anonymous):

ok thank you very much phi

OpenStudy (phi):

yw